OpenStudy (ohmath123):

what is the units digits of 943^29

OpenStudy (misssmartiez):

943 x itself by 29 times is what it is asking.

OpenStudy (misssmartiez):

Do you wish for it in expanded form?...

OpenStudy (welshfella):

Does this make sense? 943^2 last digit 9 943^3 7 943^4 1 943^5 3 So 943^9 last digit 3 can we continue this to power 29?

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} & \dfrac{943^{29}}{10} \hspace{.33em}\\~\\ & \implies \dfrac{3^{29}}{10} \hspace{.33em}\\~\\ & \implies \dfrac{3\cdot 3^{28}}{10} \hspace{.33em}\\~\\ & \implies \dfrac{3\cdot 9^{14}}{10} \hspace{.33em}\\~\\ & \implies \dfrac{3\cdot (-1)^{14}}{10} \hspace{.33em}\\~\\ & \implies \dfrac{3\cdot 1}{10} \hspace{.33em}\\~\\ & \implies 3 \hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (welshfella):

yes My way would aaalso give 3

OpenStudy (abdullahm):

@mathmath333 I understand @welshfella 's method, but I'm trying to understand your method. Why did you replace 9 with -1? Also, why did you change 3^(28) to 9^(14)?

OpenStudy (welshfella):

the last part ; 3^28 = (3^2)^14 = 9^14

OpenStudy (mathmath333):

In terms of modular arithmetic when 9 is divided by 10 the remainder is either 9 or -1.

OpenStudy (abdullahm):

Welshfella, I already knew that. I was just wondering if there was a specific reason to change the base. Mathmath33, that makes sense! Thanks to both of you!